How do you solve this equation?
Susan and Laura are asked to write an expression about a mystery number. Susan writes five less than one-third of the mystery number. Laura writes the sum of negative three and one-sixth of the mystery number. If Susan and Laura have the same mystery number, which of the following can be used to find x, the mystery number?

Answer by stanbon(75887)

How do you solve this equation?
Susan and Laura are asked to write an expression about a mystery number. Susan writes five less than one-third of the mystery number. Laura writes the sum of negative three and one-sixth of the mystery number. If Susan and Laura have the same mystery number, which of the following can be used to find x, the mystery number?
-------
Let the number be "x"..
---
Susan writes (1/3)x-5
Laura writes (1/6)x-3
----------------
Equation:
(1/3)x-5 = (1/6)x-3
(2/6)x-5 = (1/6)x-3
(1/6)x = 2
x = 12
------------
Cheers,
Stan H.
------------